OpenStudy (anonymous):

What cosine function represents an amplitude of 2, a period of 2π, a horizontal shift of π, and a vertical shift of −1?

OpenStudy (anonymous):

f(x) = −1 cos πx + 2 f(x) = −1 cos (x − π) + 2 f(x) = 2 cos (x − π) − 1 f(x) = 2 cos πx − 1

OpenStudy (anonymous):

@ash2326

OpenStudy (anonymous):

\[y=a\cos (bx-c)+d=a\cos\left(b\left(x-\frac{c}{b}\right)\right)+d\] \(|a|\) is the amplitude, \(\dfrac{2\pi}{b}\) is the period, \(\dfrac{c}{b}\) is the horizontal shift, and \(d\) is the vertical shift.

OpenStudy (anonymous):

You're given that \(|a|=2\), which means \(a=2\) or \(a=-2\). You can eliminate the first two choices.

OpenStudy (anonymous):

The period is \(2\pi\), which means \[\frac{2\pi}{b}=2\pi~~\iff~~b=1\] So you can eliminate the fourth choice.

OpenStudy (anonymous):

omg thank you, can you explain the rest please?

OpenStudy (anonymous):

Also, can you help me a few more? Trig- is not my strongest subject

OpenStudy (anonymous):

@SithsAndGiggles

OpenStudy (anonymous):

Explain the rest? What else is there to explain?

OpenStudy (anonymous):

you explain how to find the amplitude and the period

OpenStudy (anonymous):

how do find the horizontal and vertical shifts

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Okay, so \(d\) is pretty easy to identify. The vertical shift is -1 (or 1 unit down), so \(d=-1\).

OpenStudy (anonymous):

You have a horizontal shift of \(\pi\), so \[\frac{c}{b}=\pi\] We figured out that \(b=1\), so \(c=\pi\).

OpenStudy (anonymous):

Thank you. You're amazing. Can you help me with another?

OpenStudy (anonymous):

Sure, but post it as a different question. I might not respond in a timely fashion, I'm a bit busy here, so someone else may be able to assist you.

OpenStudy (anonymous):

Ok, can I still tag your name?

OpenStudy (anonymous):

I'm closing the question :)

OpenStudy (anonymous):

Wait so what one is it?